OpenStudy (anonymous):

Question 3. Faye is selling floral arrangements. Each arrangement uses 1 vase and 14 daisies. Each vase costs Faye $3.00. Let C be the total cost of the arrangement and d be the cost of 1 daisy. Which equation should Faye use to find the total cost of each arrangement? (Points : 1) C = 3d + 14 C = 14d + 3 14 = C +2d 14C = d + 3

OpenStudy (anonymous):

i think it would be B. C=14d+3 am i right?

OpenStudy (asib1214):

C=14+3

OpenStudy (asib1214):

C=14d+3

OpenStudy (asib1214):

so lets say if the total arrangement cost about $500 you will replace 500 with C and solve . 500=14d+3 and if the number of d changes you simple plug-in whatever is d..lets say 25 you'll solve for d. 500=14(25) +3 500=350d+3 500-3=350d 497=350d.....divide both sides by 350 and get your d. 1.42=d i hope i helped you.